# Multi-Agent Guided Policy Optimization

### v1

@@ 6 Conclusion
We presented MAGPO, a novel framework that bridges the gap between CTCE and CTDE in cooperative MARL. MAGPO leverages a sequentially executed guider for coordinated exploration while constraining it to remain close to the decentralized learner policies. This design enables stable and effective guidance without sacrificing deployability. Our approach builds upon the principles of GPO and introduces a practical training algorithm with provable monotonic improvement. Empirical results across 43 tasks in 6 diverse environments demonstrate that MAGPO consistently outperforms state-of-the-art CTDE methods and is competitive with CTCE methods, despite relying on decentralized execution. Looking ahead, we believe MAGPO offers a strong foundation for future MARL research. One promising direction is to exploit privileged state information during training to further enhance the guider, as in single-agent GPO. Another is to explore adaptive mechanisms for tuning the guider-learner alignment based on observed imitation gaps. We hope MAGPO contributes to more robust, scalable, and deployable multi-agent systems in real-world applications.
@@ Appendix A Proofs
Following the derivation from Xiao [45], the PMD objective is
@@
Then, by the performance difference lemma [17], we obtain:
@@
We first decompose the guider policy in (14)
@@ Appendix B CoordSum Details
[…]
where <num_agents> is the number of agents in the team, and <num_actions> specifies the size of each agent’s discrete action space.
@@
At each timestep $t\in[1,100]$, all agents receive the same observation: the current target value $target[t]\in[0,M]$, where $M$ is the maximum possible target sum (defaulting to <num_actions> - 1). The observation also includes the current step count.
@@
[…]
@@ C.1 Per-task sample efficiency results
C.2 Per-task tabular results
In table 1, we provide absolute episode metric [8, 13] over training averaged across 10 seeds with std reported. The bolded value means the best performance across all methods, while highlighted value represents the among CTDE methods.
Table 1: Performance comparison across tasks. Best overall value is bolded. Best among CTDE methods is highlighted.
Task MAGPO CTDS IPPO MAPPO MAT SABLE
CoordSum CoordSum-3x10 $107.97\pm 0.04$ $90.74\pm 5.16$ $107.08\pm 0.39$ $107.44\pm 0.34$ $92.03\pm 8.02$ $\boldmath\mathbf{133.04\pm 4.02}$
CoordSum-3x30 $118.67\pm 3.48$ $69.48\pm 5.60$ $114.20\pm 5.89$ $113.31\pm 4.80$ $50.87\pm 4.33$ $\boldmath\mathbf{132.44\pm 5.91}$
CoordSum-5x20 $117.06\pm 0.44$ $61.01\pm 5.83$ $113.84\pm 1.73$ $114.58\pm 0.47$ $67.06\pm 8.22$ $\boldmath\mathbf{131.75\pm 7.04}$
CoordSum-8x15 $\boldmath\mathbf{111.23\pm 3.15}$ $46.14\pm 19.30$ $81.70\pm 40.58$ $87.57\pm 35.57$ $60.83\pm 6.57$ $101.85\pm 24.21$
LevelBasedForaging 15x15-3p-5f $\boldmath\mathbf{0.99\pm 0.00}$ $0.96\pm 0.02$ $0.90\pm 0.03$ $0.97\pm 0.02$ $0.91\pm 0.02$ $0.96\pm 0.01$
15x15-4p-3f $\boldmath\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
15x15-4p-5f $\boldmath\mathbf{0.99\pm 0.00}$ $0.99\pm 0.00$ $0.97\pm 0.00$ $0.98\pm 0.01$ $0.97\pm 0.01$ $0.99\pm 0.00$
2s-8x8-2p-2f-coop $1.00\pm 0.00$ $1.00\pm 0.00$ $\boldmath\mathbf{1.00\pm 0.00}$ $\boldmath\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $\boldmath\mathbf{1.00\pm 0.00}$
2s-10x10-3p-3f $0.97\pm 0.01$ $0.87\pm 0.02$ $0.99\pm 0.00$ $\boldmath\mathbf{1.00\pm 0.00}$ $0.97\pm 0.01$ $0.99\pm 0.00$
8x8-2p-2f-coop $\boldmath\mathbf{1.00\pm 0.00}$ $\boldmath\mathbf{1.00\pm 0.00}$ $\boldmath\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
10x10-3p-3f $1.00\pm 0.00$ $1.00\pm 0.00$ $\boldmath\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $0.99\pm 0.00$ $1.00\pm 0.00$
MaConnector con-5x5x3a $\boldmath\mathbf{0.94\pm 0.01}$ $0.93\pm 0.02$ $0.88\pm 0.02$ $0.87\pm 0.02$ $0.85\pm 0.02$ $0.92\pm 0.02$
con-7x7x5a $\boldmath\mathbf{0.76\pm 0.03}$ $0.71\pm 0.05$ $0.65\pm 0.03$ $0.63\pm 0.02$ $0.62\pm 0.04$ $0.74\pm 0.03$
con-10x10x10a $\boldmath\mathbf{0.42\pm 0.03}$ $0.37\pm 0.04$ $0.33\pm 0.02$ $0.30\pm 0.01$ $0.22\pm 0.05$ $0.39\pm 0.03$
con-15x15x23a $0.02\pm 0.01$ $0.02\pm 0.01$ $0.05\pm 0.01$ $0.02\pm 0.01$ $00.00\pm 0$ 0.00 $\boldmath\mathbf{0.08\pm 0.01}$
MPE simple_spread_3ag $-6.10\pm 0.21$ $-6.34\pm 0.29$ $-8.35\pm 0.84$ $-6.72\pm 0.22$ $-6.59\pm 0.23$ $\boldmath\mathbf{-4.92\pm 0.30}$
simple_spread_5ag $-18.67\pm 0.41$ $-20.38\pm 0.39$ $-21.97\pm 0.47$ $-22.84\pm 0.23$ $-25.30\pm 1.74$ $\boldmath\mathbf{-12.75\pm 0.91}$
simple_spread_10ag $-40.51\pm 0.80$ $-40.09\pm 0.73$ $-42.08\pm 0.45$ $-41.83\pm 0.52$ $-50.07\pm 1.72$ $\boldmath\mathbf{-36.93\pm 0.32}$
RobotWarehouse large-4ag $\boldmath\mathbf{7.63\pm 0.98}$ $5.00\pm 0.54$ $1.84\pm 1.84$ $3.02\pm 2.26$ $4.61\pm 0.25$ $6.22\pm 1.73$
large-4ag-hard $\boldmath\mathbf{4.56\pm 0.64}$ $2.25\pm 1.50$ $0.05\pm 0.06$ $0.00\pm 0.01$ $2.28\pm 1.88$ $3.46\pm 1.80$
large-8ag $10.40\pm 0.52$ $7.68\pm 0.69$ $4.87\pm 2.48$ $8.35\pm 0.66$ $\boldmath\mathbf{14.72\pm 0.79}$ $11.01\pm 0.51$
large-8ag-hard $8.66\pm 0.61$ $4.32\pm 2.86$ $3.63\pm 2.43$ $3.38\pm 3.10$ $9.07\pm 0.71$ $\boldmath\mathbf{9.22\pm 0.48}$
medium-4ag $11.46\pm 1.16$ $8.22\pm 0.56$ $2.58\pm 1.40$ $7.82\pm 3.24$ $7.62\pm 3.83$ $\boldmath\mathbf{12.74\pm 1.44}$
medium-4ag-hard $\boldmath\mathbf{8.49\pm 0.54}$ $4.81\pm 0.79$ $1.89\pm 1.90$ $2.80\pm 2.77$ $4.64\pm 2.54$ $6.79\pm 1.34$
medium-6ag $\boldmath\mathbf{14.78\pm 3.28}$ $8.49\pm 1.07$ $3.47\pm 2.93$ $12.13\pm 0.53$ $13.32\pm 0.63$ $12.97\pm 1.03$
small-4ag $15.09\pm 0.71$ $11.07\pm 0.81$ $3.69\pm 4.03$ $10.52\pm 0.75$ $\boldmath\mathbf{18.27\pm 0.53}$ $16.47\pm 8.26$
small-4ag-hard $11.48\pm 0.41$ $7.56\pm 1.02$ $2.27\pm 2.73$ $9.44\pm 0.35$ $9.68\pm 3.19$ $\boldmath\mathbf{12.02\pm 1.20}$
tiny-2ag $19.70\pm 1.16$ $13.70\pm 1.96$ $7.61\pm 6.29$ $12.28\pm 5.86$ $17.06\pm 1.61$ $\boldmath\mathbf{21.17\pm 1.24}$
tiny-2ag-hard $\boldmath\mathbf{16.61\pm 0.99}$ $9.23\pm 0.88$ $6.15\pm 5.71$ $13.60\pm 0.73$ $13.44\pm 2.41$ $15.93\pm 0.74$
tiny-4ag $31.02\pm 1.95$ $14.42\pm 1.16$ $16.98\pm 5.06$ $26.29\pm 2.88$ $28.19\pm 1.02$ $\boldmath\mathbf{43.56\pm 2.69}$
tiny-4ag-hard $20.49\pm 2.61$ $11.31\pm 0.88$ $9.06\pm 8.18$ $19.01\pm 1.31$ $20.54\pm 11.99$ $\boldmath\mathbf{30.97\pm 1.65}$
xlarge-4ag $\boldmath\mathbf{5.74\pm 0.71}$ $3.02\pm 1.25$ $0.01\pm 0.01$ $3.73\pm 1.19$ $4.71\pm 0.43$ $3.76\pm 2.30$
xlarge-4ag-hard $0.31\pm 0.64$ $0.12\pm 0.34$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.39\pm 1.01$ $\boldmath\mathbf{0.70\pm 1.39}$
Smax 2s3z $\boldmath\mathbf{1.00\pm 0.00}$ $0.99\pm 0.01$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.00$ $1.00\pm 0.00$
3s5z $0.99\pm 0.01$ $0.98\pm 0.01$ $\boldmath\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.01$
3s5z_vs_3s6z $\boldmath\mathbf{0.95\pm 0.02}$ $0.89\pm 0.04$ $0.90\pm 0.04$ $0.91\pm 0.05$ $0.93\pm 0.03$ $0.94\pm 0.02$
3s_vs_5z $0.99\pm 0.01$ $0.97\pm 0.01$ $\boldmath\mathbf{0.99\pm 0.01}$ $0.99\pm 0.01$ $0.96\pm 0.01$ $0.99\pm 0.01$
6h_vs_8z $\boldmath\mathbf{1.00\pm 0.00}$ $0.99\pm 0.01$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
10m_vs_11m $\boldmath\mathbf{0.98\pm 0.02}$ $0.80\pm 0.21$ $0.91\pm 0.02$ $0.39\pm 0.07$ $0.88\pm 0.19$ $0.83\pm 0.24$
5m_vs_6m $0.34\pm 0.35$ $0.15\pm 0.23$ $\boldmath\mathbf{0.92\pm 0.01}$ $0.28\pm 0.37$ $0.68\pm 0.36$ $0.59\pm 0.41$
smacv2_5_units $\boldmath\mathbf{0.83\pm 0.02}$ $0.80\pm 0.03$ $0.76\pm 0.02$ $0.76\pm 0.02$ $0.82\pm 0.02$ $0.78\pm 0.03$
smacv2_10_units $0.76\pm 0.05$ $0.65\pm 0.06$ $0.65\pm 0.02$ $\boldmath\mathbf{0.76\pm 0.02}$ $0.71\pm 0.04$ $0.65\pm 0.04$
27m_vs_30m $0.99\pm 0.01$ $0.99\pm 0.01$ $0.92\pm 0.10$ $0.83\pm 0.10$ $0.88\pm 0.09$ $\boldmath\mathbf{1.00\pm 0.00}$
C.3 Hyperparameters
The default hyperparameters for all methods are listed in Table 2 and Table 3. The full hyperparameter search spaces are provided in Table 4, Table 5, Table 6, and Table 7. All algorithms are tuned on each task using the Tree-structured Parzen Estimator (TPE) implemented in the Optuna library [2], with a tuning budget of 40 trials per task.
For all environments except CoordSum, baseline hyperparameters are directly adopted from the Sable paper [23], which already provides optimized configurations under the same budget. For MAGPO, we use the same base hyperparameters as the adopted CTCE method, and only tune the additional parameters specific to MAGPO as shown in Table 7. For the CoordSum environment, full hyperparameter tuning is conducted for all algorithms.
@@ C.3 Hyperparameters
Table 3: Default hyperparameters for IPPO and MAPPO.
@@
Number transformer heads $\{1,2,4\}$
Number transformer blocks $\{1,2,3\}$
@@
Table 6: Hyperparameter Search Space for IPPO and MAPPO.

### v2

@@ 6 Conclusion
We presented MAGPO, a novel framework that bridges the gap between CTCE and CTDE in cooperative MARL. MAGPO leverages a sequentially executed guider for coordinated exploration while constraining it to remain close to the decentralized learner policies. This design enables stable and effective guidance without sacrificing deployability. Our approach builds upon the principles of GPO and introduces a practical training algorithm with provable monotonic improvement. Empirical results across 43 tasks in 6 diverse environments demonstrate that MAGPO consistently outperforms state-of-the-art CTDE methods and is competitive with CTCE methods, despite relying on decentralized execution.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China [grant numbers U22A2062, U23B2037, 12272008, 62450001, 62476008].
@@ Appendix A Proofs
We will consider policy mirror descent (PMD) objective (Shani et al., 2019):
$\mu^{(k+1)}=\underset{\mu}{\arg\max}\left\{\eta_{k}\langle\nabla V_{\rho}(\mu^{(k)}),\mu\rangle-\frac{1}{1-\gamma}D_{d_{\rho}(\mu^{(k)})}(\mu,\mu^{(k)})\right\},$ (12)
where $\eta_{k}$ is the step size, $\rho\in\Delta(\mathcal{S})$ is an arbitrary state distribution and $d_{\rho}(\mu^{(k)})$ is the discounted state-visitation distribution under the policy $\mu^{(k)}$, $D_{d_{\rho}(\mu^{(k)})}$ is the weighted Bregman divergence. Considering that
$\langle\nabla V_{\rho}(\mu^{(k)}),\mu\rangle=\sum_{s\in\mathcal{S}}\langle\nabla_{s}V_{\rho}(\mu^{(k)}),\mu(\cdot|s)\rangle,$ (13)
we obtain
$\mu^{(k+1)}$ $=\underset{\mu}{\arg\max}\left\{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}(\mu^{(k)})(\eta_{k}\langle Q^{\mu^{(k)}}(s,\cdot),\mu(\cdot|s)\rangle-D(\mu(\cdot|s),\mu^{(k)}(\cdot|s)))\right\},$ (14)
$=\underset{\mu}{\arg\max}\left\{\sum_{s\in\mathcal{S}}(\eta_{k}\langle Q^{\mu^{(k)}}(s,\cdot),\mu(\cdot|s)\rangle-D(\mu(\cdot|s),\mu^{(k)}(\cdot|s)))\right\}.$
In this paper, we use KL divergence as a special case of Bregman divergence.
@@
Following the derivation from Xiao (2022), the PMD objective is
@@
Then, by the performance difference lemma (Kakade & Langford, 2002), we obtain:
@@
We first decompose the guider policy in equation 17
@@ Appendix B CoordSum Details
$\texttt{CoordSum-<num\_agents>}\times\texttt{<num\_actions>}-\texttt{<max\_target>}$
where <num_agents> is the number of agents in the team, <num_actions> specifies the size of each agent’s discrete action space, and <max_target> is the maximum possible target sum.
@@
At each timestep $t\in[1,100]$, all agents receive the same observation: the current target value $target[t]\sim U(0,\texttt{<max\_target>})$. The observation also includes the current step count.
@@
$\mathcal{A}_{i}=\{0,1,\dots,\texttt{<num\_actions>}-1\}$
for $i=1,\dots,\texttt{<num\_agents>}$. The joint action is the vector of all agents’ selected integers.
@@ C.1 Per-task sample efficiency results
C.2 Per-task tabular results
In Table 1, we provide absolute episode metric (Colas et al., 2018; Gorsane et al., 2022) over training averaged across 10 seeds with std reported. The bolded value means the best performance across all methods, while highlighted value represents the best among CTDE methods.
Table 1: Performance comparison across tasks. Best overall value is bolded. Best among CTDE methods are underlined.
Task MAGPO CTDS HAPPO MAPPO MAT SABLE
CoordSum 3x10-30 $53.13\pm 1.41 11$$11.98\pm 11.15$ $153.32\pm 1.89$ $\mathbf{156.37\pm 3.56}$ $68.29\pm 16.80$ $153.70\pm 3.77$
3x30-50 $156.62\pm 1.59$ $76.27\pm 14.10$ $129.35\pm 5.22$ $\mathbf{158.05\pm 4.40}$ $87.79\pm 8.79$ $125.04\pm 7.18$
5x20-80 $\mathbf{157.61\pm 3.89}$ $9.62\pm 11.79 11$$19.19\pm 10.05$ $142.51\pm 4.87$ $86.86\pm 8.01$ $48.35\pm 15.76$
8x15-100 $\mathbf{129.94\pm 8.47}$ $23.96\pm 15.92$ $77.60\pm 5.09$ $127.32\pm 12.76$ $57.22\pm 4.97$ $28.91\pm 18.75$
LevelBasedForaging 15x15-3p-5f $\mathbf{0.99\pm 0.00}$ $0.96\pm 0.02$ $0.91\pm 0.02$ $0.97\pm 0.02$ $0.91\pm 0.02$ $0.96\pm 0.01$
15x15-4p-3f $\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
15x15-4p-5f $\mathbf{0.99\pm 0.00}$ $0.99\pm 0.00$ $0.89\pm 0.02$ $0.98\pm 0.01$ $7\pm 0.01 0.99±0$$9\pm 0.00$
2s-8x8-2p-2f-coop $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $\mathbf{1.00\pm 0.00}$
2s-10x10-3p-3f $0.97\pm 0.01$ $0.87\pm 0.02$ $0.99\pm 0.01$ $\mathbf{1.00\pm 0.00}$ $7\pm 0.01 0.99±0$$9\pm 0.00$
8x8-2p-2f-coop $\mathbf{1.00\pm 0.00}$ $\mathbf{1.00\pm 0.00}$ $\mathbf{1.00\pm 0.00}$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
10x10-3p-3f $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $\mathbf{1.00\pm 0.00}$ $0.99\pm 0.00$ $1.00\pm 0.00$
MaConnector con-5x5x3a $\mathbf{0.94\pm 0.01}$ $0.93\pm 0.02$ $0.93\pm 0.02$ $0.87\pm 0.02$ $0.85\pm 0.02$ $0.92\pm 0.02$
con-7x7x5a $\mathbf{0.76\pm 0.03}$ $0.71\pm 0.05$ $0.67\pm 0.02$ $0.63\pm 0.02$ $0.62\pm 0.04$ $0.74\pm 0.03$
con-10x10x10a $\mathbf{0.42\pm 0.03}$ $0.37\pm 0.04$ $0.21\pm 0.03$ $0.30\pm 0.01$ $0.22\pm 0.05$ $0.39\pm 0.03$
[…]
MPE spread_3ag $6.10\pm 0.21$ $34\pm 0.29 6.6$$63\pm 0.49$ $6.72\pm 0.22$ $6.59\pm 0.23$ $\mathbf{-4.92\pm 0.30}$
spread_5ag $18.67\pm 0.41$ $20.38\pm 0.39$ $3.42\pm 0.53 22$$2.84\pm 0.23$ $25.30\pm 1.74$ $\mathbf{-12.75\pm 0.91}$
spread_10ag $40.51\pm 0.80$ $40.09\pm 0.73$ $43.68\pm 0.55$ $41.83\pm 0.52$ $50.07\pm 1.72$ $\mathbf{-36.93\pm 0.32}$
RobotWarehouse large-4ag $\mathbf{7.63\pm 0.98}$ $5.00\pm 0.54$ $0.61\pm 0.54$ $3.02\pm 2.26$ $4.61\pm 0.25$ $6.22\pm 1.73$
large-4ag-hard $\mathbf{4.56\pm 0.64}$ $2.25\pm 1.50$ $0\pm 0.00 0.00±0$$0\pm 0.01$ $2.28\pm 1.88$ $3.46\pm 1.80$
large-8ag $10.40\pm 0.52$ $7.68\pm 0.69$ $0.00\pm 0.00$ $8.35\pm 0.66$ $\mathbf{14.72\pm 0.79}$ $11.01\pm 0.51$
large-8ag-hard $8.66\pm 0.61$ $4.32\pm 2.86$ $0.00\pm 0.00$ $3.38\pm 3.10$ $9.07\pm 0.71$ $\mathbf{9.22\pm 0.48}$
medium-4ag $11.46\pm 1.16$ $8.22\pm 0.56$ $4.04\pm 0.63$ $7.82\pm 3.24$ $7.62\pm 3.83$ $\mathbf{12.74\pm 1.44}$
medium-4ag-hard $\mathbf{8.49\pm 0.54}$ $4.81\pm 0.79$ $3.75\pm 0.63$ $2.80\pm 2.77$ $4.64\pm 2.54$ $6.79\pm 1.34$
medium-6ag $\mathbf{14.78\pm 3.28}$ $8.49\pm 1.07$ $4.99\pm 0.82$ $12.13\pm 0.53$ $13.32\pm 0.63$ $12.97\pm 1.03$
small-4ag $5.09\pm 0.71 11$$1.07\pm 0.81$ $9.64\pm 2.43$ $10.52\pm 0.75$ $\mathbf{18.27\pm 0.53}$ $16.47\pm 8.26$
small-4ag-hard $11.48\pm 0.41$ $7.56\pm 1.02$ $6.50\pm 1.09$ $9.44\pm 0.35$ $9.68\pm 3.19$ $\mathbf{12.02\pm 1.20}$
tiny-2ag $19.70\pm 1.16$ $13.70\pm 1.96$ $10.93\pm 2.30$ $12.28\pm 5.86$ $17.06\pm 1.61$ $\mathbf{21.17\pm 1.24}$
tiny-2ag-hard $\mathbf{16.61\pm 0.99}$ $9.23\pm 0.88$ $9.61\pm 3.09$ $13.60\pm 0.73$ $13.44\pm 2.41$ $15.93\pm 0.74$
tiny-4ag $31.02\pm 1.95$ $14.42\pm 1.16$ $17.84\pm 2.17$ $26.29\pm 2.88$ $28.19\pm 1.02$ $\mathbf{43.56\pm 2.69}$
tiny-4ag-hard $20.49\pm 2.61$ $11.31\pm 0.88$ $16.66\pm 2.50$ $19.01\pm 1.31$ $20.54\pm 11.99$ $\mathbf{30.97\pm 1.65}$
xlarge-4ag $\mathbf{5.74\pm 0.71}$ $3.02\pm 1.25$ $0.00\pm 0.00$ $3.73\pm 1.19$ $4.71\pm 0.43$ $3.76\pm 2.30$
xlarge-4ag-hard $0.31\pm 0.64$ $12\pm 0.34 0.0$$\pm 0.00 0.00±0$$0\pm 0.00$ $0.39\pm 1.01$ $\mathbf{0.70\pm 1.39}$
Smax 2s3z $\mathbf{1.00\pm 0.00}$ $9\pm 0.01 0.99±0$$9\pm 0.01$ $1.00\pm 0.00$ $0.99\pm 0.00$ $1.00\pm 0.00$
3s5z $0.99\pm 0.01$ $0.98\pm 0.01$ $0.98\pm 0.01$ $1.00\pm 0.00$ $\mathbf{1.00\pm 0.00}$ $1.00\pm 0.01$
3s5z_vs_3s6z $\mathbf{0.95\pm 0.02}$ $0.89\pm 0.04$ $0.51\pm 0.11$ $0.91\pm 0.05$ $0.93\pm 0.03$ $0.94\pm 0.02$
[…]
6h_vs_8z $\mathbf{1.00\pm 0.00}$ $9\pm 0.01 0.99±0$$9\pm 0.00$ $1.00\pm 0.00$ $0.99\pm 0.01$ $1.00\pm 0.00$
10m_vs_11m $\mathbf{0.98\pm 0.02}$ $0.80\pm 0.21$ $0.14\pm 0.03$ $0.39\pm 0.07$ $0.88\pm 0.19$ $0.83\pm 0.24$
5m_vs_6m $0.34\pm 0.35$ $15\pm 0.23 0.0$$03\pm 0.01$ $0.28\pm 0.37$ $\mathbf{0.68\pm 0.36}$ $0.59\pm 0.41$
smacv2_5_units $\mathbf{0.83\pm 0.02}$ $0.80\pm 0.03$ $0.68\pm 0.02$ $0.76\pm 0.02$ $0.82\pm 0.02$ $0.78\pm 0.03$
smacv2_10_units $0.76\pm 0.05$ $0.65\pm 0.06$ $0.47\pm 0.06$ $\mathbf{0.76\pm 0.02}$ $0.71\pm 0.04$ $0.65\pm 0.04$
27m_vs_30m $9\pm 0.01 0.99±0$$9\pm 0.01$ $0.77\pm 0.05$ $3\pm 0.10 0.88±0$$8\pm 0.09$ $\mathbf{1.00\pm 0.00}$
C.3 Hyperparameters
All algorithms were tuned on each task with a tuning budget of 40 trials using the Tree-structured Parzen Estimator (TPE) implemented in the Optuna library (Akiba et al., 2019). Since some of the tuned hyperparameters are provided in Mava (de Kock et al., 2023), we directly adopt them and only tune the additional algorithms and tasks. Specifically, MAGPO and HAPPO in all tasks, all algorithms in the newly introduced CoordSum tasks.
The default hyperparameters for all methods are listed in Table 2 and Table 3. The full hyperparameter search spaces are provided in Table 4, Table 5, Table 6, and Table 7.
@@ C.3 Hyperparameters
Table 3: Default hyperparameters for HAPPO and MAPPO.
@@
Number of transformer heads $\{1,2,4\}$
Number of transformer blocks $\{1,2,3\}$
@@
Table 6: Hyperparameter Search Space for HAPPO and MAPPO.